If a§b = a^b + b^a for all real numbers a,b, what is (2§3)§1?
Answer by user_dude2008(1862)

2§3 = 2^3 + 3^2 = 8+9 = 17

(2§3)§1 = 17§1 = 17^1 + 1^17 = 17 + 1 = 18